OpenStudy (anonymous):

Find the value of x and y so that the quadrilateral is a parallelogram. x = 8, y = 9 x = 8, y = 23.5 x = 15, y = 32 x = 15, y = 9

OpenStudy (anonymous):

4x – 17 = 2x – 1 2x – 17 = -1 2x = 16 x = 8 3y + 5 = 5y – 13 5 = 2y – 13 18 = 2y 9 = y